OpenStudy (anonymous):

Maureen deposited $19,000 at 11 percent interest compounded annually. (a) Writean equation that expresses the amount of money in the account as a function of n, where n equals the number of years since the original deposit was made. (b) Use the equation in (a) to determine how much money is in the account at the end of n = 8 years. (c) What was the total interest earned at n = 8 years?

OpenStudy (anonymous):

got anything?

OpenStudy (anonymous):

rudi got anything?

OpenStudy (anonymous):

(a) A = P(1 + i)^n A - is the final amount P - is the initial amount(money deposited) i - is the interest rate(always written in decimal form) n - is the number of years am not sure. Is this what you looking for?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

its good enough thank u

OpenStudy (anonymous):

ok that's for number (a)...let me do other questions

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Rudi..? by chance do you know how to enter that formula into graphing calc.

OpenStudy (anonymous):

At the en of 8 years. A= unknown p = $19,000 n = 8 years i = r/100 = 11/100. There for i = 0.11 From here, we will substitude the given data on our equation A = P(1+ i)^n try it..

OpenStudy (anonymous):

total interest earned?

OpenStudy (anonymous):

Total interet earned (I) = Total amount (A) - Initial amount (P) use this equation

OpenStudy (anonymous):

Did you get the answer to number (b) ?

OpenStudy (anonymous):

yeah give me the numbers for the total interest

OpenStudy (anonymous):

total interest for n = 8 years

OpenStudy (anonymous):

A= 19,000 (1+0.11)^8

OpenStudy (anonymous):

remember that all the numbers are on your question(refer back to the question) and the final amount aka total amount is the answer in (b). i.e. A = $43786.22 p = $ 19000

OpenStudy (anonymous):

I = A - p = 43786.22 - 19000 = 24786.22 This means Maureen earned a total interest of $24786.22 at the end of 8 years